Suppose a DC power source with a voltage of 45 volts is connected to a 15 Ω load. How much power will this load dissipate?
1
Expert's answer
2020-05-26T12:48:17-0400

Explanations & Calculations

  • This could easily be calculated using the relationships p = Vi & V = iR for a resistive load.

"\qquad\qquad
\begin{aligned}
\small p & = \small V \times i \,\,\, \text{and} \,\,\, i = \frac{V}{R}\\
\small p & =\small \frac{V^2}{R}\\
&= \small \frac{45^2V}{15\Omega}\\
&= \small \bold{135W}
\end{aligned}"